OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17 Question 2 Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −one halfx + 3 y = −one halfx + 6 y = 2x + 3 Question 3 Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = −one halfx − 2 y = 2x

OpenStudy (anonymous):

the prependicular line is the reciporal of the slope and the opposite sign

OpenStudy (anonymous):

I dont understand this one bit i went over it like three times and i still dont understand

OpenStudy (anonymous):

okso the equation they are using is mx + b m being slope x being the x coodinate and b being the y intercept

OpenStudy (anonymous):

ok could you help me do one so i could see how its done if you dont mind

OpenStudy (anonymous):

for a perpendicular line it is the reciprocal and the opposite sign so that will make -1/3x+ y=7

OpenStudy (anonymous):

now we are going to have to but it in slope intercept form mx+b can you do that?

OpenStudy (anonymous):

im still kind of lost sorry if im being trouble

OpenStudy (anonymous):

ya lets start over...

OpenStudy (anonymous):

first we know that the lines pass through (6, −1). right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok wel lets use this formula (y - y1) = m(x - x1)

OpenStudy (texaschic101):

do ya'll need help ?

OpenStudy (anonymous):

ya i need to eat dinner so can you take it from here @texaschic101

OpenStudy (texaschic101):

I got you covered...go eat

OpenStudy (anonymous):

ok and if you would like to help i need all the help i could get thanks

OpenStudy (texaschic101):

if you are looking for a perpendicular line, the slope will be the negative reciprocal. Your equation is 3x + y = 7. We want it in y = mx + b form where m is the slope. So lets put it in y = mx + b. 3x + y = 7 y = -3x + 7 are you with me so far ?

OpenStudy (anonymous):

ya im with you

OpenStudy (texaschic101):

ok....the slope for the equation given is -3....do you understand how I got that ?

OpenStudy (anonymous):

not sure really

OpenStudy (texaschic101):

we put you equation in y = mx + b and m is the slope. y = -3x + 7.....in this form, the slope is the number right before the x. So the slope for this equation is -3. Understand ?

OpenStudy (anonymous):

ok now i understand

OpenStudy (texaschic101):

for a perpendicular line, the slope is the negative reciprocal....what that means is...well..let me just show you....-3 is the same as -3/1....what we do is flip the number and change the sign....making the slope we want 1/3. Do you understand that ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

then we use this formula : y - y1 = m(x - x1) using slope(m) = 1/3 and points ( 6,-1), y1 = -1 and x1 = 6 and we sub all that in to the formula understand ?

OpenStudy (anonymous):

yep

OpenStudy (texaschic101):

y - y1 = m(x - x1) y - (-1) = 1/3(x - 6) y + 1 = 1/3x - 2 y = 1/3x - 2 - 1 y = 1/3x - 3 Are you understanding what I have done ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

Now when you are looking for perpendicular line....it is the negative reciprocal ...flip the number and change the sign. But when you are looking for a parallel line, the slope does not change. You have to first get the given equation in y = mx + b form, and the slope is m. So, you want to try one ?

OpenStudy (anonymous):

i would like to but my work is being time and i really have to get it done but thanks for your help it was very helpful

OpenStudy (texaschic101):

you are very welcome :)